. NEED ANSWER ASAP / ANSWER NEVER USED BEFORE, COMPLETELY NEW ANSWER PLEASE

Current and Quick Ratios

The Nelson Company has $1,287,000 in current assets and $495,000 in current liabilities. Its initial inventory level is $335,000, and it will raise funds as additional notes payable and use them to increase inventory.

a.)How much can Nelson's short-term debt (notes payable) increase without pushing its current ratio below 2.2? Do not round intermediate calculations. Round your answer to the nearest dollar. $  

b.)What will be the firm's quick ratio after Nelson has raised the maximum amount of short-term funds? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

a) When the company raise funds as additional notes payable and use them to increase inventory, the inventories and notes payable of the company will increases, which lead to increase in current assets and current liabilities.

Suppose Nelson can increase its short-term debt without pushing its current ratio= X

Current ratio= Current assets/Current liabilities

2.20= ($1287000+X)/($495000+X)

$1287000+X= 2.20($495000+X)

$1287000+X= 1089000+2.20X

X= $165000

Nelson's short-term debt (notes payable) increase without pushing its current ratio below 2.2= X= $165000

b) New current assets= $1287000+165000= $1452000

New inventories= $335000+165000= $500000

Quick assets= Current assets-Inventories

= $1452000-500000= $952000

New current liabilities= $495000+165000= $660000

Quick ratio= Quick assets/Current liabilities

= $952000/660000= 1.44